Supplementary Operations Report – April 12th

MDRS Supplemental Operations Report 12-APR-2021

Name of person filing report: Shannon Rupert

Reason for filing report: Follow up to beginning of field season

Non-nominal systems: The Work Party that did not happen

Action taken for non-nominal systems: Once again we would just like to thank each and every person who has come to volunteer at MDRS during a work party or refit. I have always respected, appreciated and been damn grateful for your help, and knew we couldn’t do it without you and this year showed us that is true. I am looking forward to the end of the pandemic and your return.

Generator check, note if oil and coolant added: rental generator on site and connected to our system, new generator has arrived in SLC and we will pick it up in the next two weeks.

ScienceDome AC unit: off

Solar— Functioning nominally

Diesel Reading – 50 %

Propane Reading – Outpost 1: 40 %, Outpost 2: 65%, Main Campus 75%

Ethanol Free Gasoline – 15 gallons

Water (Outpost tank) – 70 gallons

Water (static tank) – 100 gallons

Water (GreenHab) – 0 gallons

Water (loft) – 10

Static to Loft Pump used – yes

Water Meter: not noted

Toilet tank emptied: yes, twice

Perseverance rover used: yes

Hours: 199.7

Beginning charge: 100 %

Ending charge: not noted

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes, in town

Spirit rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes, in town

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: no, in town

Curiosity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes, in town

Notes and action taken on rovers: Nothing to report

Monthly check and fill of rover batteries (first workday of the month): nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) yes. The three 350’s have been prepped for crew use. Installed a new battery in 350.3.

Reason for use: crew use

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used: 1

Notes and action taken on ATVs: Installed a new battery in 350.3.

HabCar running? If used, why, where? Yes, taken to Grand Junction to pick up rental generator today

CrewCar running? If used why, where? Used yesterday, taken to Grand Junction for the crew’s use

Van (Luna) running? If used why, where? Yes, taken to Grand Junction yesterday for staff to deliver CrewCar to crew.

Campus wide inspection, if action taken, what and why? The crew completed the tunnel work today

Evidence of rodents, where and action taken: Nothing new to report.

Summary of internet: All nominal.

Summary of suits and radios: To be reported by Crew Engineer in standard Operations Report

Summary of Hab operations: Water tank cleaning completed. Noted one knob (far left) on the stove has been broken. Burner is still usable but a replacement knob should be ordered. A water leak was noticed in the line from the static tank to the loft tank and needs to be replaced before next season. Remote power system monitoring is back online and nominal.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We still need to clean this building but quite honestly, I am exhausted and only have one good hand so I will be finishing it after the crew goes into sim.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: All crew tested negative for COVID prior to arrival. Masks and social distancing are being used consistently and proactively by the crew. My thanks to them.

List and summary of projects: Nothing to report at this time.

Questions, concerns, supplies needed and requests: Welcome back to our Capcoms and Mission Support personnel. We have missed you!